When three pumps, A, B, and C, are running together, they can pump 3500 gal per hour. When only A and B are running, 2100 gal per hour can be pumped. When only A and C are running, 2400 gal per hour can be pumped. What is the pumping capacity of each pump?
Answer by Alan3354(21608)

When three pumps, A, B, and C, are running together, they can pump 3500 gal per hour. When only A and B are running, 2100 gal per hour can be pumped. When only A and C are running, 2400 gal per hour can be pumped. What is the pumping capacity of each pump?
--------------
A + B + C = 3500
A + B = 2100
A + C = 2400
---------------
A + B + C = 3500
A + B = 2100
---------------- Subtract
C = 1400
A = 1000
B = 1100